# You Only Sample Once: Taming One-Step Text-To-Image Synthesis by Self-Cooperative Diffusion GANs

### v2

@@ Abstract
We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis.This is achieved by integrating the diffusion process with GANs. Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning. In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without explicit training.
@@ 1 Introduction
Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [41, 58, 6, 7], text-to-video [4, 13], image editing [9, 5, 33] and controlled generation [61, 34]. However, the generation process of DMs requires iterative denoising, leading to slow generation speed. This exacerbates the computational demands of large-scale DMs. The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
@@
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50]. Some previous works [47, 28, 29, 2] focus on developing advanced ODE-solver, for reducing the sampling steps. However, they still require typically 20+ steps to achieve high-quality generation. Another line is to distill models from pre-trained PF-ODEs [49, 27, 30, 31, 42], aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods [30, 31] can achieve reasonable sample quality with 4 to 8 steps. However, it is still challenging to generate high-quality samples with one step.
@@
In contrast, Generative Adversarial Networks (GANs) [8, 39] are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [43, 17]. In this work, we present a novel approach to combine diffusion process and GANs. Previous works [52, 57, 58] have developed some variants of diffusion-GAN hybrid models. However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation. In order to directly learn the one-step generation by adversarial divergence and avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process. And we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence. This approach can be regarded as a self-cooperative process, as the generator learns from itself. We find the innovative design enables stable training and effective learning for one-step generation. Hence, we named our model YOSO which means You Only Sample Once. Besides, we extend our introduced algorithm to one-step text-to-image generation based on pre-trained models by several principled designs. Thanks to our effective designs, we can fine-tune existing text-to-image pre-trained models (i.e., Stable Diffusion [41] and PixArt- [6]) for high-quality one-step generation (See Figure 1). It is worth noting that we even can develop one-step text-to-image models by Low Rank Adaptation (LoRA) [16] fine-tuning.
@@ 2 Related Works
Few-step Text-to-image Generation. Recently, several works have been developed for few-step text-to-image generation. The LCM [30] builds upon consistency distillation [49], adapting it to Stable Diffusion (SD) and enabling image generation in 4 steps with acceptable quality. Nevertheless, reducing the sampling steps led to poor generation quality. InstaFlow [27] builds upon the Rectified Flows [26, 25], adapting it to Stable Diffusion, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. Moreover, due to the model’s training on fixed data-noise pairings, it loses the flexibility to support variable data-noise combinations, which is detrimental to tasks like image modification or iterative multi-step sampling that can enhance the quality of generated images. Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen extends DDGANs [52] and SSIDMs [57] to Stable Diffusion by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD [44] introduces a one-step text-to-image generation based on stable diffusion. It follows earlier research by employing a pre-trained image encoder, DINOv2 [35], as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands.Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth. Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to stable diffusion but also to PixArt- [6] which is based on Diffusion Transformer [37]. This demonstrates the wide application of our proposed YOSO.
@@ 3 Background
Diffusion Models. Diffusion models [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12]. A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN Hybrids. An issue in diffusion models is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs [52] propose to minimize the adversarial divergence between model and :
where and is imposed by a GAN generator. The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps) compared to Gaussian.
@@ 4.1 Revisiting of Diffusion GANs
Diffusion-GAN hybrid models are tailored for training with a large denoising step size. However, a key issue is that they match the generator distribution at corrupted data distribution. This formulation only indirectly learns the and , which are the distribution used for one-step generation, making the learning process less effective.
@@ 4.2 Our Design
The optimization objective is constructed by the combination of adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
@@
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Models [49]. However, Consistency Models consider the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE related to denoising step size.
@@ 5 Towards One-Step Text-to-Image Synthesis
Since training a text-to-image model from scratch is quite expensive, we suggest using pre-trained text-to-image diffusion models as initialization with Self-Coopeartive Diffusion GANs. In this section, we introduce several principled designs for developing a generation model that enables one-step text-to-image synthesis based on pre-trained DMs.
@@ 5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies [27, 49] has found that the LPIPS loss [62] is crucial for obtaining few-step diffusion models with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular Latent diffusion models operate in the latent space to reduce computational demands. Hence, using LPIPS loss in training Latent diffusion models is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained Stable Diffusion can serve as an effective feature extractor [56], we suggest using pre-trained Stable Diffusion to perform latent perceptual loss. However, Stable Diffusion is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
where is the latent encoded images by VAE and is the text feature. We note that the benefit of computing latent perceptual loss by stable diffusion is not only the computational efficiency but also the incorporation of text features, which are crucial in text-to-image tasks.
Latent Discriminator. Training GANs for the text-to-image task on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained text-to-image diffusion has learned representative features. Hence, we suggest to apply the pre-trained stable diffusion for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a more stable training process with fast convergence. Note that, expensive R1 regularization is not needed in our approach.
@@ 5.2 Fixing the Noise Scheduler
A common issue in diffusion models is that the final corrupted samples are not pure noise. For example, the noise schedule used by the Stable Diffusion makes terminal Signal-to-noise ratio (SNR) far from reaching zero. The corrupted samples at the final timestep follows:
@@
where the terminal SNR is , effectively creates a gap between training and inference. Previous work [24] has only observed that such an issue makes diffusion models unable to produce pure black or white images, which does not hurt sample quality seriously. However, we find that this issue yields serious problems in one-step generation. As shown in Figure 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21]. Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below:
@@
where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. The proposed Informative prior initialization effectively addresses the "prior hole issue" in one-step generation. As shown in Figure 2, the artifacts in one-step generation are immediately removed. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRa fine-tuning.
@@
Quick Adaption to V-prediction and Zero Terminal SNR. The Informative prior initialization is not a perfect solution, and it suffers from numerical instability when terminal SNR is pretty low. As shown in Figure 3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. As discussed in [24], the solution should be the combination of v-prediction [42] and zero terminal SNR. However, we found that directly switching to v-prediction and zero terminal SNR is computationally expensive, requiring lots of extra iterations. This is unacceptable in solving large-scale text-to-image with limited computational resource. To address this, we propose an adapt-stage to fine-tune pre-trained diffusion models to v-prediction and zero terminal SNR.
@@
Adapt-stage-I switch to v-prediction : where , , denotes the parameters of the desired v-prediction model, denotes the parameters of the frozen per-trained model with -prediction.
@@
Adapt-stage-II switch to zero terminal SNR : Note that in adapt-stage-II, we change the noise scheduler to zero terminal SNR. But for the teacher models, we keep the noise scheduler as the same. This approach avoids numerical instability issues of -prediction with zero terminal SNR. We note that the zero terminal SNR has lower SNR than the original schedule at each timestep, so the loss formulates an effective objective to construct a zero terminal SNR student by distilling from the teacher.
@@
We notice that this adapt-stage converges rapidly, typically requiring only 1k iterations for training. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction models, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise schedule.
@@ 6 Experiments
In this section, we evaluate the proposed YOSO for the image synthesis problem. In Section 6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies to highlight the effectiveness of our algorithm. Section 6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [41]. In Section 6.4, we show that the YOSO can be used for several downstream applications.
@@ 6.1 Unconditional Image Generation
Experiment Setting. For the model architectures, we use the same UNet and discriminator architecture as used in DDGANs [52]. We set the denoising steps as four. For the evaluation metric, we choose the Fréchet inception distance (FID) [10]. We train the proposed model on the CIFAR-10 dataset [60]. Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
Results. The quantitative results are shown in Table 1. We observe that our method provides competitive performance compared with the best DMs and GANs. Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours. We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation. Note that we do not employ data augmentation in experiments. We also observe that the consistency models built upon the diffusion process provide competitive and even better results compared to us. However, we note that the training of consistency models requires tons of computational resources and careful design of the time schedule. As shown in Figure 4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations. To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours. Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO. We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.

### v3

@@ Abstract
We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis. YOSO integrates the diffusion process with GANs to achieve the best of two worlds. Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning. In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training.
@@ 1 Introduction
Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [40, 60, 6, 7], text-to-video [4, 13], image editing [9, 5, 32] and controlled generation [64, 33]. However, the generation process of DMs requires iterative denoising, leading to slow generation speed. The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
@@
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50]. Some previous works [47, 27, 28, 2] focus on developing advanced ODE-solver, for reducing the sampling steps. However, they still require 20+ steps to achieve high-quality generation. Another line is distilling from pre-trained PF-ODEs [49, 26, 29, 30, 41], aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods [29, 30] can achieve reasonable sample quality with 4+ steps. However, it is still challenging to generate high-quality samples with one step.
@@
In contrast, generative adversarial networks (GANs) [8, 38] are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [42, 17]. In this work, we present a novel approach to combine diffusion process and GANs. Previous works [54, 59, 60] have developed some variants of diffusion-GAN hybrid models. However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation. To directly learn the one-step generation by adversarial divergence and to avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process, where denotes the distribution of noisy data and denotes the denoised distribution parameterized by denoising generator. Then, we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence. This approach can be regarded as a self-cooperative process, as the generator learns from itself. Such an innovative design enables stable training and effective learning for one-step generation. Hence, we name our model YOSO representing You Only Sample Once. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler).
Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion [40] and PixArt- [6]) for high-quality one-step generation (See Fig. 1). Moreover, we are the first to support Low Rank Adaptation (LoRA) [16] fine-tuning for one-step generation of text-to-image.
@@ 2 Related Works
Recently, several works have been developed for few-step text-to-image generation. LCM [29] adapts consistency distillation [49] for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow [26] adapts rectified flows [25, 24] for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. DMD [62] adapts variation score distillation [52] for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen [60] extends DDGANs [54] and SSIDMs [59] to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD [44] introduces a one-step text-to-image generation based on SD. It follows earlier research [43] by employing a pre-trained image encoder, DINOv2 [34], as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth. Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt- [6] which is based on diffusion transformer [36]. This demonstrates the wide application of our proposed YOSO.
@@ 3 Background
Diffusion Models. Diffusion models (DMs) [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12]. A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN Hybrids. An issue in DMs is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs [54] propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator. The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps).
@@ 4.1 Revisiting of Diffusion GANs
A key issue in Diffusion-GAN hybrid models is that they match the generator distribution at corrupted data distribution. This formulation only indirectly learns the and , which are the distributions used for one-step generation, making the learning process less effective.
@@ 4.2 Our Design
where and . The optimization objective is constructed by the combination of adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
@@
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Training [49]. However, Consistency Training considers the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE.
@@ 5 Towards One-Step Text-to-Image Synthesis
Since training a text-to-image model from scratch is quite expensive, we suggest using pre-trained text-to-image DMs as initialization with Self-Coopeartive Diffusion GANs. In this section, we introduce several principled designs for developing a generation model that enables one-step text-to-image synthesis based on pre-trained DMs.
@@ 5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies [26, 49] has found that the LPIPS loss [65] is crucial for obtaining few-step DMs with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular SD operate in the latent space to reduce computational demands. Hence, using LPIPS loss in training latent DMs is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained SD can serve as an effective feature extractor [58], we suggest using pre-trained SD to perform latent perceptual loss. However, SD is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
where is the latent encoded images by VAE and is the text feature. We note that the benefit of computing latent perceptual loss by SD is not only the computational efficiency but also the incorporation of text features, which are crucial in text-to-image tasks.
Latent Discriminator. Training GANs for the text-to-image on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained SD has learned representative features. Hence, we suggest applying the pre-trained SD for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work [44] that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a stable training process with fast convergence. Note that, expensive R1 regularization is not needed in our approach.
Decoupled Scheduler. We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical. Hence, to better unleash the capability of pre-trained DMs, we propose using decoupled schedulers to construct the following loss in fine-tuning pre-trained DMs:
where , , and we let and in experiments.
@@ 5.2 Fixing the Noise Scheduler
A common issue in DMs is that the final corrupted samples are not pure noise. For example, the noise scheduler used by the SD makes the corrupted samples at the final timestep as: , where the terminal Signal-to-noise ratio (SNR) is , effectively creates a gap between training and inference. Previous work [23] has only observed that this makes DMs unable to produce pure black or white images. However, we find that this issue yields serious problems in one-step generation. As shown in Fig. 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
@@
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21]. Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below: , where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. As shown in Fig. 2, the artifacts in one-step generation are immediately removed after applying IPI. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRA fine-tuning.
@@
Quick Adaption to V-prediction and Zero Terminal SNR. The IPI suffers from numerical instability when terminal SNR is pretty low. As shown in Fig. 3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. Following [23], we suggest switching to v-prediction [41] and zero terminal SNR. However, we found that direct transition convergences slowly (see Appendix F). This is unacceptable in solving large-scale text-to-image with limited computational resources. To address this, we propose a quick adapt-stage:
@@
Adapt-stage-I switch to v-prediction : where , , denotes the desired v-prediction model, and denotes the frozen pre-trained model.
@@
Adapt-stage-II switch to zero terminal SNR : Note that in this stage, we only change the scheduler to zero terminal SNR for student model. This avoids numerical instability issues of -prediction with zero terminal SNR. We note that the zero terminal SNR scheduler has lower SNR than the original schedule at each timestep, formulating an effective distillation objective.
@@
We note that this adapt-stage converges rapidly, typically requiring only 1k iterations for training. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction DMs, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise scheduler.
@@ 6 Experiments
In this section, we evaluate the performance of the proposed YOSO. In Sec. 6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies in Sec. 6.2 to highlight the effectiveness of our algorithm. Sec. 6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [40]. In Sec. 6.4, we show that the YOSO can be used for several downstream applications.
@@ 6.1 Unconditional Image Generation
Experiment Setting. We use the same UNet and discriminator architecture as used in DDGANs [54]. We set the denoising steps as four. For the evaluation metric, we choose the Fréchet inception distance (FID) [10]. We train the proposed model on the CIFAR-10 dataset [63]. Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
Results. The quantitative results are shown in Tab. 1. We observe that our method provides competitive performance compared with the best DMs and GANs. Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours. We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation. We also observe that the consistency models provide competitive and even better results compared to us. However, we note that the training of consistency models requires tons of computational resources and careful design of the noise schedule. As shown in Fig. 4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations. To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours. Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO. We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.